OpenStudy (anonymous):

Solve the equation. Be sure to show your work. All final answers must be in simplest form. x+7/11=9/11 i need a little assistance

OpenStudy (anonymous):

x = 2/11...

OpenStudy (anonymous):

x+7/11=9/11 x=9/11-7/11 You have common denominator

OpenStudy (anonymous):

so you can solve it or just plug in the answer that alf gave

OpenStudy (anonymous):

so i did have the idea of the 2 but i did not know that i should put it in a fraction form thanks

OpenStudy (anonymous):

Yes it is a better idea to put it in a fraction form.

OpenStudy (anonymous):

thanks for the help it has been so long since i did this type of math